# Only a Little to the Left: A Theory-grounded Measure of Political Bias in Large Language Models

### v1

@@ Abstract
Prompt-based language models like GPT4 and LLaMa have been used for a wide variety of use cases such as simulating agents, searching for information, or for content analysis. For all of these applications and others, political biases in these models can affect their performance. Several researchers have attempted to study political bias in language models using evaluation suites based on surveys, such as the Political Compass Test (PCT), often finding a particular leaning favored by these models. However, there is some variation in the exact prompting techniques, leading to diverging findings and most research relies on constrained-answer settings to extract model responses. Moreover, the Political Compass Test is not a scientifically valid survey instrument. In this work, we contribute a political bias measured informed by political science theory, building on survey design principles to test a wide variety of input prompts, while taking into account prompt sensitivity. We then prompt 11 different open and commercial models, differentiating between instruction-tuned and non-instruction-tuned models, and automatically classify their political stances from 88,110 responses. Leveraging this dataset, we compute political bias profiles across different prompt variations and find that while PCT exaggerates bias in certain models like GPT3.5, measures of political bias are often unstable, but generally more left-leaning for instruction-tuned models.
@@
Mats Faulborn1, Indira Sen2, Max Pellert3, Andreas Spitz1, and David Garcia1 1University of Konstanz 2University of Mannheim 3Barcelona Super Computing Center Correspondence: David.Garcia@uni-konstanz.de
@@ 1 Introduction
In the past years, Large Language Models (LLMs) have emerged as a transformative technology with applications in a large variety of social domains including medicine Singhal et al. (2023); Thirunavukarasu et al. (2023); Zhou et al. (2024), finance Jeong (2024); Li et al. (2023c); Wu et al. (2023), education Elkins et al. (2023); Kasneci et al. (2023) and academia Beltagy et al. (2019); Meyer et al. (2023); Porsdam Mann et al. (2023). More and more people use these technologies for complex information integration tasks such as search to become informed about or to summarise historical and current events Sharma et al. (2024). However, researchers and policy-makers have found evidence of societal bias in language technologies, including LLMs. Political bias is especially problematic in some of these information gathering, search, or even content analysis-related tasks, because they can perpetuate certain existing real-life biases Sharma et al. (2024). Given these challenges, researchers, policy-makers, industry and other involved stakeholders must ensure that emerging technologies do not contribute to such problems but rather, in an optimistic view, help to solve them. To achieve this, it is of vital importance to develop valid measures of political bias in LLMs.
@@
Careful bias analyses require evaluating varied inputs and outputs of the models with evaluation techniques that have high construct validity. However, most research on political bias in LLMs is limited to static settings where the model is forced to answer with a single answer token, thereby only evaluating highly constrained outputs, that are not ecologically valid, i.e., in real-world usage, users of LLMs rarely restrict them to single tokens. So far, few studies have investigated political bias in an open-answer setting, while also accounting for prompt sensitivity Feng et al. (2023); Röttger et al. (2024); Wright et al. (2024). Adding to these problems, no clear conceptual definition of political bias has been brought forward and most previous work relies on poorly documented and popular science survey inventories like the Political Compass Test (PCT). In addition, a large portion of the literature focuses on investigating closed-source models that lack openness about the technical details of the system and stringent documentation of possible model changes that go undocumented Fujimoto and Takemoto (2023); Hartmann et al. (2023); Motoki et al. (2023); Rozado (2023); Rutinowski et al. (2024).
@@
This study addresses these gaps by 1) introducing a theory-driven definition of political bias that relies on a scientifically valid survey inventory — the political leaning items in the World Values Survey EVS/WVS (2022), 2) developing a political bias measure with inherent consideration of 30 prompt variations applied over eleven commercial and open-weight models. Using this novel methodology, we find that instruction-tuned models exhibit considerable left-leaning political bias. However, we also show that PCT exaggerates political bias for certain models, the variation of input prompts significantly affects the resulting bias measures, and that constrained answer settings lead to unpredictable model outputs. Based on our finding, we provide concrete suggestions for measuring political bias in LLMs. Our code, data, and prompts are available.111anonymized for review
@@ 2 Operationalizing Political Bias of Language Models
While there has been a plethora of recent work on political bias in LLMs (see Section 5.1), there is still a lack of a general definition of political bias. One notable exception is Liu et al. (2022), who distinguish between political bias that is measured when a political entity is mentioned in the prompt (direct bias) or not mentioned in the prompt (indirect bias). However, this definition still does not explain the concept of political bias itself. Generally speaking, when talking about political bias, we mean divergences of political attitudes and ideas on a ideological spectrum from left to right. In political science, ideology is generally defined as a set of political ideas that are interconnected and stable Campbell et al. (1960), as well as interdependent Converse (1964). In the context of LLMs, this implies that there needs to be some consistency to the political ideas fabricated by a model. In practice this implies that if a model argues for both sides of the political spectrum, we cannot reliably position it on an ideological scale and therefore the model would not be considered politically biased or ideological. Recent political science literature suggests that ideology among the US public is better described with a left-right scale along two dimensions rather than one Carmines et al. (2012b). Therefore, we disaggregate our political bias measure into one cultural and one economic dimension.
Toward a Political Bias Measure. Given language model responses to political statements and political leaning extracted from these responses (see Section 3.4), we define the political bias measures as follows. For a model $m$ and political directions $d\in\left\{left,right\right\}$ we define the count of model answers agreeing with statements $d$ as $A$, disagreeing with $d$ as $D$ and neutral answers w.r.t $d$ as $N$. Agreement of $m$ with $d$ is now computed as the proportion of answers that agree with $d$ relative to all other valid answers not labelled as unrelated:
$P_{agree,m,d}=\frac{A}{A+D+N}$ (1)
@@
Drawbacks of the Political Compass Test. To create a corpus of political statements, we rely on the Political Compass Test or PCT The Political Compass (2023) and on parts of the European Values Study and World Values Survey (WVS) Joint Questionnaire EVS/WVS (2022), which enables us to collect a total of 89 political statements. While the PCT has been widely used in recent work to measure political bias in LLMs Feng et al. (2023); Röttger et al. (2024), it is not a scientific instrument. Unlike WVS, it has little documentation on how it was developed, how the propositions were pretested, and with whom. All of this information is crucial to gauge the validity of survey instruments Pitt et al. (2021). Furthermore, the PCT has several irrelevant propositions such as “Astrology accurately explains many things.” and loaded propositions (c.f., Table 3), which specifically steer responses and are discouraged by survey methodologists Clark and Schober (1992).
In contrast, the World Values Survey has been widely used by researchers as well as survey institutes to measure sociocultural attitudes, including political leaning, for several decades  [^2] https://www.worldvaluessurvey.org/WVSContents.jsp. By creating a corpus of political statements sourced from the WVS, we contribute a theory-grounded measure of political bias in LLMs.
@@ 3.1 Prompting Setup
We use prompts to evaluate political bias that are made up of two parts: The prefix used to ask the question and the political statement to which the model responds. We vary both to elicit as much variance in the model answers as possible.
As a first step, we determine the political bias of possible responses to the statements in PCT and WVS by creating labels that indicate what political side is reflected by approving or disapproving each statement. We, i.e., two authors of the paper, manually label the statements according to two conditions: 1) Whether agreeing with the statement reflects the right or left side of the political spectrum and 2) Whether the statement concerns economic or cultural issues. The annotators are both fluent English speakers, one has a bachelor’s degree in political science and the other in cognitive science. Cohen’s $\kappa$ between the manual annotators is 0.77 for the first condition and 0.76 for the second.
Next, we use GPT-4 to reformulate the statement in two additional ways, resulting in three different versions of the same statement. 1) For the first version, we prompt GPT-4 to reformulate the original statement by changing its wording but retaining the meaning, since we want to test whether the responses of the model are influenced by simple rewording of the statement. 2) The second reformulation prompts GPT-4 to reverse the statement to reflect the other side of the political spectrum. The rationale for this is to test whether the evaluated language model is willing to reflect on the other side of the political spectrum with respect to one specific statement. Following the definition of political bias brought forward in this work, if a model agrees with both formulations of the statement, we would not consider it as being politically biased. Table 1 provides one example to illustrate the reformulation process.
@@
To ensure the consistency of the reformulations, one of the authors of the paper manually checked all of them. In two cases out of 178, the model failed once to reword the original statement and once to reverse the political meaning. To solve these problematic cases, it was enough to make GPT-4 aware of the error and to pose the task again. For the political opposite condition, we flip the labels ‘agree’ and ‘disagree’ to align the meaning of agreeing or disagreeing with a statement over all three conditions. The prompts used for creating and labeling the statements are in Appendix A. As a further robustness check and because we use GPT-4 for creating the dataset but the model is also among the models tested for bias, we reran our analysis with Llama-3.1-405B-Instruct. The difference in results are negligible (c.f Appendix H).
@@ 3.2 Prefix Variants
Röttger et al. (2024) have shown that using different prefixes when prompting models to answer the Political Compass Test significantly alters model responses and consequently also impacts the positioning on the resulting scale. Therefore, in addition to the three versions of political statements, we also make use of a prefix that asks for a model response. Table 2 lists all ten prefixes, while Appendix E provides more details about each prefix.
@@
With three different versions of each of the political statements and ten different prefixes, our prompting framework queries models in 30 different ways to ensure a comprehensive evaluation of political bias, accounting for prompt sensitivity.
@@ 3.3 Language Models
Leveraging our bias audit framework, we prompt the following open-weight and commercial generative LLMs: falcon-7b, falcon-7b-instruct, falcon-40b-instruct, llama-7b-hf, llama-7b-chat-hf, llama-13b-hf, llama-13b-chat-hf, mistral-7b-v0.1, mistral-7b-instruct-v0.1, gpt-3.5-turbo-0125 and gpt-4333For querying model responses for bias computation, we use version gpt-4-0613 which was newly released at the time. Almazrouei et al. (2023); Jiang et al. (2023); Touvron et al. (2023). The models are chosen because they are widely used and represent four model families that are build by different model creators, possibly influencing the level of political bias generated by these models. For every LLM and prompt combination, we obtain three answers or runs. The final dataset has 88,110 observations (30 prompt versions * 3 runs * 11 models * 89 statements).
@@ 3.4 Measuring Political Stance from LLM Answers
For all but one prefix, we prompt our models to generate free-text answers, from which political stance labels need to be extracted. We use an additional model to detect the stance of responses. We first employed the approach by Feng et al. (2023) that uses zero-shot inference, leveraging BART-Large Lewis et al. (2019), which is fine-tuned on the Multilingual Natural Language Inference (MNLI) dataset Williams et al. (2018). However, when validating the zero-shot classifier on a manually annotated test set, our results (Figure 3) showed that the zero-shot capabilities of BART-Large do not yield satisfactory performance on our data. Domain specific fine-tuning, on the other hand, has been shown to be an effective strategy for enhancing the capabilities of zero-shot classifiers Chae and Davidson (2023). Therefore, we collect a stratified random sample from the data to fine-tune the classifier. [^4] All hyperparameters are reported in Appendix D. We randomly select four observations from each prompt-model pair, totaling a training dataset of 1320 unique model answers, comprising about 1.5% of the original data. Following best practices in text annotation for automatic classifier development Mendelsohn et al. (2021); Barberá et al. (2021), this training dataset of 1320 instances is single-coded by one of the authors, while the test set (described below) is consensus coded, i.e., annotated by multiple annotators. The labels assigned are: {agree, disagree, neutral, unrelated}.
Validating the Political Stance Classifier. To ensure that the stance classifier is reliable, we compare it’s performance against manual annotations. Two authors of the paper independently annotate 264 randomly sampled instances of LLM answers for political stance. The two annotator have substantial agreement (Cohen’s $\kappa=$0.68) and disagreements are resolved in a discussion round. We then compare the performance of the automatic stance classifier against these manual annotations.
We evaluate the performance on two metrics: a macro-averaged F1-Score and the number of observations maintained when excluding observations that are predicted with a certain level of confidence. The confidence corresponds to the raw entailment probabilities. The F1-Score of the classifier approximately doubled to a 0.93 macro averaged F1-score when looking at a 0.9 confidence level while retaining about 5 times more data. To ensure accurately labelled data, we exclude all observations with a lower confidence than 0.9 leading to a loss of about 33% of the data in the final dataset. [^5] More details about the performance of the stance classifier can be found in Appendix C.
@@
Table 3: GPT3.5’s response to two items from the Political Compass Test and World Values Survey on the topic of Government assistance. The framing of the statements lead to opposite responses.
Figure 1: Political bias for instruction vs. base models over the two dimensions of political ideology, disaggregated by the measurement instruments, i.e., World Values Survey (filled markers) and Political Compass Test (unfilled markers).

### v2

@@ Abstract
Prompt-based language models like GPT4 and LLaMa have been used for a wide variety of use cases such as simulating agents, searching for information, or for content analysis. For all of these applications and others, political biases in these models can affect their performance. Several researchers have attempted to study political bias in language models using evaluation suites based on surveys, such as the Political Compass Test (PCT), often finding a particular leaning favored by these models. However, there is some variation in the exact prompting techniques, leading to diverging findings, and most research relies on constrained-answer settings to extract model responses. Moreover, the Political Compass Test is not a scientifically valid survey instrument. In this work, we contribute a political bias measured informed by political science theory, building on survey design principles to test a wide variety of input prompts, while taking into account prompt sensitivity. We then prompt 11 different open and commercial models, differentiating between instruction-tuned and non-instruction-tuned models, and automatically classify their political stances from 88,110 responses. Leveraging this dataset, we compute political bias profiles across different prompt variations and find that while PCT exaggerates bias in certain models like GPT3.5, measures of political bias are often unstable, but generally more left-leaning for instruction-tuned models. Code and data are available on GitHub [^1] https://github.com/MaFa211/theory_grounded_pol_bias.
@@
Mats Faulborn1, Indira Sen2, Max Pellert3, Andreas Spitz4, and David Garcia4thanks: Corresponding Author 1scieneers GmbH 2University of Mannheim 3Barcelona Super Computing Center 4University of Konstanz mats.faulborn@gmx.net, indira.sen@uni-mannheim.de, max.pellert@bsc.es, andreas.spitz@uni-konstanz.de, david.garcia@uni-konstanz.de
@@ 1 Introduction
In the past years, Large Language Models (LLMs) have emerged as a transformative technology with applications in a large variety of social domains including medicine Singhal et al. (2023); Thirunavukarasu et al. (2023); Zhou et al. (2024), finance Jeong (2024); Li et al. (2023c); Wu et al. (2023), education Elkins et al. (2023); Kasneci et al. (2023) and academia Beltagy et al. (2019); Meyer et al. (2023); Porsdam Mann et al. (2023). More and more people use these technologies for complex information integration tasks, such as search, to become informed about or to summarise historical and current events Sharma et al. (2024). However, researchers and policy-makers have found evidence of societal bias in language technologies, including LLMs. Political bias is especially problematic in some of these information gathering, search, or even content analysis-related tasks, because they can perpetuate certain existing real-life biases Sharma et al. (2024). Given these challenges, researchers, policy-makers, industry and other involved stakeholders must ensure that emerging technologies do not contribute to such problems but rather, in an optimistic view, help to solve them. To achieve this, it is of vital importance to develop valid measures of political bias in LLMs.
@@
Careful bias analyses require evaluating varied inputs and outputs of the models with evaluation techniques that have high construct validity. However, most research on political bias in LLMs is limited to static settings where the model is forced to answer with a single answer token, thereby only evaluating highly constrained outputs that are not ecologically valid, i.e., in real-world usage, users of LLMs rarely restrict them to single tokens. So far, few studies have investigated political bias in an open-answer setting, while also accounting for prompt sensitivity Feng et al. (2023); Röttger et al. (2024); Wright et al. (2024). Adding to these problems, no clear conceptual definition of political bias has been brought forward and most previous work relies on poorly documented and popular science survey inventories like the Political Compass Test (PCT). In addition, a large portion of the literature focuses on investigating closed-source models that lack openness about the technical details of the system and stringent documentation of possible model changes that go undocumented Fujimoto and Takemoto (2023); Hartmann et al. (2023); Motoki et al. (2024); Rozado (2023); Rutinowski et al. (2024).
@@
This study addresses these gaps by 1) introducing a theory-driven definition of political bias that relies on a scientifically valid survey inventory — the political leaning items in the World Values Survey and European Values Survey EVS/WVS (2022), 2) developing a political bias measure with inherent consideration of 30 prompt variations applied over 11 commercial and open-weight models. Using this novel methodology, we find that instruction-tuned models exhibit considerable left-leaning political bias. However, we also show that the PCT exaggerates political bias for certain models, that the variation of input prompts significantly affects the resulting bias measures, and that constrained answer settings lead to unpredictable model outputs. Based on our findings, we provide concrete suggestions for measuring political bias in LLMs.
@@ 2 Operationalizing Political Bias of Language Models
While there has been a plethora of recent work on political bias in LLMs (see Section 5.1), there is still a lack of a general definition of political bias. One notable exception is Liu et al. (2022), who distinguish between political bias that is measured when a political entity is mentioned in the prompt (direct bias) or not mentioned in the prompt (indirect bias). However, this definition still does not explain the concept of political bias itself. Generally speaking, when talking about political bias, we mean divergences of political attitudes and ideas on a ideological spectrum from left to right. In political science, ideology is generally defined as a set of political ideas that are interconnected and stable Campbell et al. (1960), as well as interdependent Converse (1964). In the context of LLMs, this implies that there needs to be some consistency to the political ideas fabricated by a model. In practice, this implies that if a model argues for both sides of the political spectrum, we cannot reliably position it on an ideological scale and therefore the model would not be considered politically biased or ideological. Recent political science literature suggests that ideology among the US public is better described with a left-right scale along two dimensions rather than one Carmines et al. (2012b). Therefore, we disaggregate our political bias measure into one cultural and one economic dimension.
[…]
@@
Drawbacks of the Political Compass Test. To create a corpus of political statements, we rely on the Political Compass Test or PCT The Political Compass (2023) and on parts of the European Values Study (EVS) and World Values Survey (WVS) Joint Questionnaire EVS/WVS (2022), which enables us to collect a total of 89 political propositions. While the PCT has been widely used in recent work to measure political bias in LLMs Feng et al. (2023); Röttger et al. (2024), it was not developed as a questionnaire using standard social scientific methodology. The PCT uses Likert scales, asking respondents to rate their level of agreement or disagreement with each statement. There is no documentation or peer-reviewed research on how these items were developed or whether they were pretested or if so, with whom. All of this information is crucial to gauge the validity of survey instruments Pitt et al. (2021), constituting a basic level of testing needed before using a questionnaire in social scientific research. Between 2001, the year of the creation of the PCT, and 2019, only one article in the Dimensions.ai database mentions the PCT, and does it in the context of its online popularity rather than as a scientific instrument. In 2020, it is mentioned for the first time as a possibility to study political bias in word embeddings Gordon et al. (2020), leading to 71 further papers mentioning the PCT between 2021 and 2024, the vast majority mentioning it in the context of text analysis or LLM research. Up to May 2025, in Google scholar, just 143 articles mention the Political Compass Test. We reviewed all of them and found no article explaining the development of the questionnaire or its validity in any empirical study.
The website of the PCT acknowledges creating biased and loaded propositions in the introduction to the test: “To question the logic of individual ones that irritate you is to miss the point”, “Some propositions are extreme, and some are moderate” The Political Compass (2023). In addition, the PCT has irrelevant propositions, such as “Astrology accurately explains many things” and loaded propositions (c.f., Table 4), which specifically steer responses and are discouraged by survey methodologists Clark and Schober (1992). The case of the astrology item is also a sign of the focus of the PCT to English-speaking countries, where belief in Astrology can be a correlate of ideology but not a fundamental issue that can be used as an item in other countries or as a stable signal over time.
In contrast, the World Values Survey has been widely used by researchers as well as survey institutes to measure sociocultural attitudes, including political leaning, for several decades [^2] https://www.worldvaluessurvey.org/WVSContents.jsp. In particular, there is established quantitative evidence in the sociology literature showing that the WVS predicts observable behaviors and other reported attitudes across countries Schwartz (2007). The World Values Survey is mentioned in 57,337 articles in Dimensions.ai since 2001, has more than 58,000 results on Google Scholar, and has been used in recent work in the NLP literature Ramezani and Xu (2023); Giuliani et al. (2024). By creating a corpus of political statements sourced from the WVS, we contribute a theory-grounded measure of political bias in LLMs. In our work, we aim to make a comparable measurement in both the scales of the PCT and the WVS, making formulations of items as comparable as possible and using the same response formats.
@@ 3.1 Prompting Setup
We use prompts to evaluate political bias that are composed of two parts: The prefix used to ask the question and the political statement to which the model responds. We vary both to elicit as much variance in the model answers as possible.
As a first step, we determine the political bias of possible responses to the statements in PCT (N = 62) and WVS (N = 27) by creating labels that indicate which political side is reflected by approving or disapproving each statement. The labelling is done according to two conditions: 1) Whether agreeing with the statement reflects the right or left side of the political spectrum and 2) Whether the statement concerns economic or cultural issues. We use GPT-4 to decide all labels used for bias computation. As a robustness check, two authors of the paper manually labeled a sample (N = 40) for verification, an approach similar to Kim et al. (2024). The annotators are both fluent English speakers, one has a bachelor’s degree in political science and the other in cognitive science. Cohen’s $\kappa$ between the manual annotators and GPT-4 is 0.77 for the first condition and 0.76 for the second, indicating substantial agreement. Therefore, we use the GPT-4 labels as the ground truth of the propositions. Table 1 shows the counts of both label categories for the two data sources used while all propositions and their labels can be found in the Appendix (Tables 1011, and 12).
Source Issue Label
Political
Label
Count
PCT cultural left 9
PCT cultural right 31
PCT economic left 10
PCT economic right 12
WVS cultural left 4
WVS cultural right 14
WVS economic left 2
WVS economic right 7
Table 1: Distribution of political statements based on data source and label.
Next, we use GPT-4 to reformulate all statements in two additional ways, resulting in three different versions of the same statement. 1) For the first version, we prompt GPT-4 to reformulate the original statement by changing its wording but retaining the meaning, since we want to test whether the responses of the model are influenced by simple rewording of the statement. 2) The second reformulation prompts GPT-4 to reverse the statement to reflect the other side of the political spectrum. The rationale for this is to test whether the evaluated language model is willing to reflect on the other side of the political spectrum with respect to one specific statement. Following the definition of political bias brought forward in this work, if a model agrees with both formulations of the statement, we would not consider it as being politically biased. Table 2 provides one example to illustrate the reformulation process.
@@
To ensure the consistency of the reformulations, one of the authors of the paper manually checked all of them. In two cases out of 178, the model failed once to reword the original statement and once to reverse the political meaning. To solve these problematic cases, it was sufficient to make GPT-4 aware of the error and to pose the task again. For the political opposite condition, we flip the labels ‘agree’ and ‘disagree’ to align the meaning of agreeing or disagreeing with a statement over all three conditions. The prompts used for creating and labeling the statements are in Appendix A. As a further robustness check and because we use GPT-4 for creating the dataset although the model is also among the models tested for bias, we reran our analysis with Llama-3.1-405B-Instruct. The differences in results are negligible (c.f Appendix I).
@@ 3.2 Prefix Variants
Röttger et al. (2024) have shown that using different prefixes when prompting models to answer the Political Compass Test significantly alters model responses and consequently also impacts the positioning on the resulting scale. Therefore, in addition to the three versions of political statements, we also make use of a prefix that asks for a model response. Table 3 lists all ten prefixes, while Appendix F provides more details about each prefix.
@@
With three paraphrased versions of every political statement and ten distinct prefixes, our prompting framework queries the models in 30 different ways, enabling a comprehensive evaluation of political bias that accounts for prompt sensitivity.  The prefixes reflect real-world LLM usage across multiple domains.  The truth prefix emulates a fact-checking request, whereas the please and non-please variants cover polite and more terse, command-style usage.  The emotion variants reproduce the persuasive, affect-laden phrasing users bring to customer-support or social-media exchanges, and the name prefix echoes the wake-word greetings now ingrained in voice-assistant culture.
@@ 3.3 Language Models
Leveraging our bias audit framework, we prompt the following open-weight and commercial generative LLMs: falcon-7b, falcon-7b-instruct, falcon-40b-instruct, llama-7b-hf, llama-7b-chat-hf, llama-13b-hf, llama-13b-chat-hf, mistral-7b-v0.1, mistral-7b-instruct-v0.1, gpt-3.5-turbo-0125 and gpt-4333For querying model responses for bias computation, we use version gpt-4-0613 which was newly released at the time. Almazrouei et al. (2023); Jiang et al. (2023); Touvron et al. (2023). The models are chosen because they are widely used and represent four model families that are build by different model creators, possibly influencing the level of political bias generated by these models. For every LLM and prompt combination, we obtain three answers (or runs). The final dataset has 88,110 observations (30 prompt versions * 3 runs * 11 models * 89 statements).
@@ 3.4 Measuring Political Stance from LLM Answers
For all but one prefix, we prompt our models to generate free-text answers, from which political stance labels need to be extracted. We use an additional model to detect the stance of responses. We first employed the approach by Feng et al. (2023) that uses zero-shot inference, leveraging BART-Large Lewis et al. (2019), which is fine-tuned on the Multilingual Natural Language Inference (MNLI) dataset Williams et al. (2018). However, when validating the zero-shot classifier on a manually annotated test set, our results (Figure 3) showed that the zero-shot capabilities of BART-Large do not yield satisfactory performance on our data. Domain specific fine-tuning, on the other hand, has been shown to be an effective strategy for enhancing the capabilities of zero-shot classifiers Chae and Davidson (2025). Therefore, we collect a stratified random sample from the data to fine-tune the classifier [^4] All hyperparameters are reported in Appendix E. We randomly select four observations from each prompt-model pair, totaling a training dataset of 1,320 unique model answers, comprising about 1.5% of the original data. Following best practices in text annotation for automatic classifier development Mendelsohn et al. (2021); Barberá et al. (2021), this training dataset of 1,320 instances is single-coded by one of the authors, while the test set (described below) is consensus coded, i.e., annotated by multiple annotators. The labels assigned are: {agree, disagree, neutral, unrelated}.
Validating the Political Stance Classifier. To ensure that the stance classifier is reliable, we compare its performance against manual annotations. Two authors of the paper independently annotate 264 randomly sampled instances of LLM answers for political stance. The two annotators have substantial agreement (Cohen’s $\kappa=0.68$) and disagreements are resolved in a discussion round. We then compare the performance of the automatic stance classifier against these manual annotations.
We evaluate the performance on two metrics: a macro-averaged F1-Score and the number of observations maintained when excluding observations that are predicted with a certain level of confidence. The confidence corresponds to the raw entailment probabilities. The F1-score of the classifier approximately doubled to a 0.93 macro averaged F1-score when looking at a 0.9 confidence level while retaining about 5 times more data. To ensure accurately labelled data, we exclude all observations with a lower confidence than 0.9 leading to a loss of about 33% of the data in the final dataset [^5] More details about the performance of the stance classifier can be found in Appendix D.
@@
Table 4: GPT3.5’s response to two statements from the Political Compass Test and World Values Survey on the topic of Government assistance. The framing of the statements lead to opposite responses.
Figure 1: Political bias for instruction-tuned vs. base models over the two dimensions of political ideology, disaggregated by the measurement instruments, i.e., World Values Survey (filled markers) and Political Compass Test (unfilled markers).